OpenStudy (muscrat123):

NEED HELP!!MEDAL AND FAN!!! A construction engineer needs to make 25 square concrete slabs with sides of length x feet and a height x - 3 feet. If the engineer can use at most 350,000ft^3 of concrete, what should the dimensions of each slab be, to the nearest foot? a. Write an expression for the volume of all the slabs b. What should x equal, to the nearest foot?

OpenStudy (muscrat123):

@campbell_st @dan815

OpenStudy (muscrat123):

@jim_thompson5910

OpenStudy (anonymous):

-3

OpenStudy (muscrat123):

what?

OpenStudy (muscrat123):

@BRONCOBURDEN17 @jim_thompson5910

OpenStudy (muscrat123):

please help

OpenStudy (muscrat123):

@horsegirl27

OpenStudy (campbell_st):

so reading the question is the side lengths are x and the height is x. base area of a slab is A = x * x multply by the height x -3 to find the volume of 1 slab so if you make 25 slab the total volume will be 25 * volume of 1 slab hope that makes sense

OpenStudy (muscrat123):

not really :(

OpenStudy (campbell_st):

so the volume for 25 can be equated to the total volume

OpenStudy (muscrat123):

i still dont understand

jimthompson5910 (jim_thompson5910):

`square concrete slabs with sides of length x feet and a height x - 3 feet.` each slab looks like this 

OpenStudy (muscrat123):

ok

jimthompson5910 (jim_thompson5910):

what is the volume of one slab?

OpenStudy (muscrat123):

it shows that picture

OpenStudy (muscrat123):

idk the volume

jimthompson5910 (jim_thompson5910):

multiply out the dimensions

jimthompson5910 (jim_thompson5910):

volume = length*width*height

OpenStudy (muscrat123):

3x-3

OpenStudy (muscrat123):

?

jimthompson5910 (jim_thompson5910):

what is x times x equal to?

OpenStudy (muscrat123):

2x

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

x+x = 2x

jimthompson5910 (jim_thompson5910):

x*x = x^2 x^2 = x squared

OpenStudy (muscrat123):

x^2

OpenStudy (muscrat123):

x^3-3?

jimthompson5910 (jim_thompson5910):

so we have x^2 times (x-3) multiply the x^2 by x to get x^3 multiply the x^2 by -3 to get -3x^2

jimthompson5910 (jim_thompson5910):

so x^2*(x-3) = x^3 - 3x^2

jimthompson5910 (jim_thompson5910):

`x^3 - 3x^2` is the volume of ONE slab but we have 25 of them all together

OpenStudy (muscrat123):

would that be the expression for part a ?

OpenStudy (muscrat123):

a. Write an expression for the volume of all the slabs b. What should x equal, to the nearest foot?

jimthompson5910 (jim_thompson5910):

it's close to the answer, but not quite there. Carefully re-read my last post

OpenStudy (muscrat123):

oh okay. x^3 - 3x^2 * 25?

jimthompson5910 (jim_thompson5910):

25 times the whole group, so you use parenthesis

jimthompson5910 (jim_thompson5910):

25*(x^3 - 3x^2) distribute to get 25*(x^3 - 3x^2) = 25x^3 - 75x^2

jimthompson5910 (jim_thompson5910):

`25x^3 - 75x^2` is the volume of 25 slabs

OpenStudy (muscrat123):

ok is that the final expression?

jimthompson5910 (jim_thompson5910):

for the 25 slabs, yes that's the total volume

OpenStudy (muscrat123):

okay now how do i solve?

OpenStudy (muscrat123):

?

jimthompson5910 (jim_thompson5910):

set 25x^3 - 75x^2 equal to 35,000 and solve 25x^3 - 75x^2 = 35,000 25x^2(x - 3) = 35,000 x^2(x - 3) = 35,000/25 x^2(x - 3) = 1,400 x^3 - 3x^2 = 1,400 x^3 - 3x^2 - 1,400 = 0 At this point, use a graphing calculator to solve for x

OpenStudy (muscrat123):

i dont understand. can u go step by step?

jimthompson5910 (jim_thompson5910):

you type in x^3 - 3x^2 - 1,400 into a calculator at this point. Do you have one?

OpenStudy (muscrat123):

yea hang on

OpenStudy (muscrat123):

-1400?

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have?

OpenStudy (muscrat123):

i have a ti-30xa so i used the website http://www.meta-calculator.com/online/?panel-201-calculator

jimthompson5910 (jim_thompson5910):

ok desmos is probably a better grapher https://www.desmos.com/calculator

jimthompson5910 (jim_thompson5910):

type in `x^3 - 3x^2 - 1400` and find out where it crosses the x axis

OpenStudy (muscrat123):

12.28?

jimthompson5910 (jim_thompson5910):

oh my bad, I used 35,000 when it should be 350,000 350,000/25 = 14,000

jimthompson5910 (jim_thompson5910):

type in `x^3 - 3x^2 - 14000` instead

OpenStudy (muscrat123):

25.14?

jimthompson5910 (jim_thompson5910):

yes, which rounds to x = 25 so that's the largest x can get

OpenStudy (muscrat123):

ok thanks!

jimthompson5910 (jim_thompson5910):

yw